OpenStudy (anonymous):

Explain how applying the Commutative Property of Multiplication to twenty-five times twenty-two times 4 makes it easier to simplify. Show the solution in your explanation.

OpenStudy (epoweritheta):

commutativity is changing the order of the numbers being multiplied . 25*22*4=22*25*4=22*100=2200 , note i first multiplied 25 and 4 to make the calculation simple